OpenStudy (jenniferjuice):

Find the vertex of the quadratic equation: y = 2x2 + 8x − 4. (−2, 20) (−2, −12) (2, 20) (−2, −28)

OpenStudy (anonymous):

2(x^2 + 4x - 2) = y x^2 + 4x - 2 = y/2 x^2 + 4x +4 = y/2 + 6 (x+2)^2 = y/2 + 6 (x+2)^2 - 6 = y/2 2(x + 2)^2 - 12 = y

OpenStudy (jenniferjuice):

Oh, okay, I get it now. thank you.,!

OpenStudy (cwrw238):

cydnimarie has converted your equation to vertex form

OpenStudy (anonymous):

You're welcome :)